determine if the quantities in each pair of ratios are proportional.
34 students from 8 schools and 25 students from 6 schools

Answer by checkley77(12844)

34/8=25/6 CROSS MULTIPLY
8*25=34*6
200=204 THESE RATIOS ARE NOT EQUAL.